If you have a gas at 300K and 1atm of pressure, and then you drop the temperature to 225K, what will the new pressure of the gas be?
1
Expert's answer
2020-06-05T05:13:08-0400

p1/T1=p2/T2;

p2=p1*T2/T1=1*225/300=0.75 atm;

Answer: 0.75 atm